Equation 𝑓(𝑧, 𝑝, π‘ž) = 0 , it’s solution is given by


1
Expert's answer
2021-07-27T06:13:37-0400
"f(z, p, q)=0"

Let "z=g(x+ay)" be a trial solution of the equation "f(z, p, q)=0." Β 

We, now put "x+ay=u," so that, we have


"z=g(u), \dfrac{\partial u}{\partial x}=1, \dfrac{\partial u}{\partial y}=a"

"p=\dfrac{\partial z}{\partial x}=\dfrac{dz}{du}\cdot\dfrac{\partial u}{\partial x}=\dfrac{dz}{du}"

"q=\dfrac{\partial z}{\partial y}=\dfrac{dz}{du}\cdot\dfrac{\partial u}{\partial y}=a\dfrac{dz}{du}"

Substituting the values of "p" and "q" in "f(z, p, q)=0," we get


"f(z, \dfrac{dz}{du}, a\dfrac{dz}{du})=0"

which is an ordinary differential equation of order one.


The solution of the equation


"f(z, \dfrac{dz}{du}, a\dfrac{dz}{du})=0"

is given by "z=g(u+b)" or "z=g(ax+y+b)," which is the complete integral of partial differential equation


"f(z, p, q)=0"